# IFT-UAM/CSIC-25-85

### v1

@@ I Introduction
The temperature of the universe has been decreasing ever since the end of inflation, when the Standard Model (SM) was reheated at its highest temperatures. For all the standard evolution of our Universe, the temperature has been decreasing relentlessly, possibly crossing the critical values, $T_{c}$, of some phase transitions. Were these phase transition of first order, there would be a potentially detectable signal in the form of a stochastic gravitational waves (GWs) background with a characteristic peak frequency, see Caprini:2015zlo ; Hindmarsh:2013xza ; Hindmarsh:2015qta ; Hindmarsh:2017gnf ; Cutting:2018tjt ; Cutting:2019zws ; Gould:2021dpm ; Weir:2017wfa ; Hindmarsh:2020hop ; Caprini:2019egz ; Badger:2022nwo ; Caldwell:2022qsj . Unfortunately, in the SM, neither the electroweak symmetry breaking nor QCD confinement are first order phase transitions. Nonetheless, these are targets of present interferometers LIGOScientific:2019vic and PTAs NANOGrav:2023gor ; EPTA:2023sfo as well as many future observatories such as LISA LISA and ET Sathyaprakash:2012jk . This is due to the fact that it is conceivable that new physics can, for example, affect the shape of the Higgs potential, making the electroweak symmetry breaking a first order phase transition Grojean:2004xa ; Delaunay:2007wb ; Ellis:2018mja ; Ellis:2019oqb ; Bruggisser:2018mus ; Espinosa:2007qk ; Beniwal:2017eik ; Barger:2007im ; Espinosa:2011ax ; Kozaczuk:2019pet ; Kurup:2017dzf ; Azatov:2022tii (see Athron:2023xlk for a review). More in general, other phase transitions can be present associated to the breaking of some symmetries, each with clear experimental targets at GW interferometers.
@@
As emphasised in Buen-Abad:2023hex , however, studying epochs where the temperature grows - while the Universe expands little - can reveal interesting new effects. Indeed, if the temperature grows above $T_{c}$ and then back below it, it might be well the case that a phase transition happens twice.
In this work, we explore this idea more, trying to understand what are the particle physics models that can show a phase transition when their sectors are heating up, possibly thank to an external reservoir. In most cases we will focus on the reheating phase of the Universe, but our results will apply to more generic scenarios (such as a freeze-in type of generation). Also, we want to identify which are those models that can be heated up maximally, in the search for large signals in GWs.
We identify scenarios of super-heating in sectors with a large number of light bosons in thermal equilibrium, as they have the properties to generate a persistent meta-stable minimum where these bosons are actually massive. The order parameter is a generic background field $\phi$, which, rather interestingly instead, does not need to be thermal. In these sectors, there is no a priori limit on the maximal reheating temperature, but only a practical limitation induced by the tunnelling. The first order transition then proceeds with bubble nucleation towards the new minimum where the states are massless. The balance of temperature and mass difference results in transitions that are of inverse type Barni:2024lkj ; Barni:2025mud , indeed, the plasma particles gain momentum in the true vacuum where they are relativistic Buen-Abad:2023hex . Other setups of phase transitions while heating can be found in Caprini:2011uz ; Casalderrey-Solana:2022rrn ; Kolesova:2023mno ; Dent:2024bhi .
We assess the importance of two aspects underlying super-heating and heating transitions in general. First, the sector has to have a very largeN-N number of bosons in order to populate parameter space where the effective potential for $\phi$ is under perturbative control, and the calculation of the tunnelling rate is reliable. Second, the sector has to be heated-up slowly, such as to make sure that the zero temperature minimum, where the bosons are massive, becomes a metastable minimum also at high temperature. We find that efficient thermal contact with the external reservoir at temperatures below the critical value during the heating phase is essential. In this regime, the system can undergo a pairs of phase transitions—an inverse transition upon heating followed by a direct transition upon cooling—and we quantify the resulting characteristic double-peaked gravitational-wave signal.
@@ II Super-heating: generalities
In this section, we sketch the basic idea behind a superheating phase transition. The aim is to highlight the needed ingredients to have a potential that has a metastable minimum that persists up to extremely high temperatures. We are mostly interested in situations where the metastability of this configuration leads to a first order phase transition via tunnelling, as this has associated production of gravitational waves from the early universe, but we will comment on more generic endpoints.
Let us consider the example of a sector with a scalar quantum field, $\phi$, at finite temperature, $T$. As it will be evident, the sector needs additional interacting light fields, but for the moment, we do not discuss the microphysics of this set-up. The typical mass scale of $\phi$ is $M$, such that at zero temperature the field has potential
@@
For $M^{2}>0$ the field breaks spontaneously a $Z_{2}$ symmetry, since $\phi=0$ is a maximum, in a minimum $\langle\phi\rangle=f$. [^1] Notice also that we are just interested in classifying the properties of super-heating, the fact of using a real scalar is just for simplicity. Therefore we are not going to discuss the domain wall generation when $Z_{2}$ is spontaneously broken. Eq. (1) is just prototypical of a sector with spontaneous symmetry breaking.
@@
If the temperature is the largest scale, and we can neglect $M$, this is also all the potential for $\phi$, and we will discuss later in section IV the impact of finite $M$. In conventional field theories, we have $a_{1}=0$, $a_{3}\leq 0$ and $a_{4}>0$, while the sign of $a_{2}$ depends on the theory Weinberg:1974hy , but it is almost always positive. Also, usually, the field is subject to the constraint $\phi\geq 0$.
@@
were the second ensures that the new minimum - that we call $\phi_{*}(T)$ from now on - is in the region $\phi\geq 0$.
Notice that crucially in this regime of ultra-high temperature, all these conditions are temperature independent. This is due to the fact that eq. (2) can be rewritten in terms of a dimensionless combination $\phi/T$, and the potential will just show an overall $T^{4}$ factor.
@@
Let us emphasise that if both (3) and (4) are satisfied, then the potential (2) has a metastable minimum away from the origin for any value of the temperature, including extremely high values. We are therefore interested in a slightly different phase than the one mainly discussed in Buen-Abad:2023hex .
If interested in a local minimum $\phi_{*}(T)$ at high temperature, it is convenient to consider models where $a_{2}$ is very large, such as to push the barrier $\phi_{b}(T)$ and the local minimum away from the origin, to improve calculability. Notice that the barrier and the local minimum always are separated by $\phi_{*}(T)-\phi_{b}(T)=(\sqrt{\Delta}/a_{4})\,T$, located at an average distance from the origin $\approx-a_{3}/(2a_{4})T$ for small $\Delta$. Taking $\Delta$ as an input, we can express $a_{4}=(a_{3}^{2}-\Delta)/4a_{2}$, which leads to $a_{3}<-3\sqrt{\Delta}$ as an implicit new form for eq. (4). For this, it is convenient to define $a_{3}=-3(1+\epsilon)\sqrt{\Delta}$, with $\epsilon>0$. Therefore we have
$\begin{split}V_{\rm min}&=T^{4}\frac{a_{2}^{3}\epsilon}{\Delta(2+3\epsilon)^{3}}\ ,\\ \frac{V_{\rm b}}{V_{\rm min}}&=\frac{1}{4\epsilon}+O(\epsilon)\ ,\quad\varphi_{\rm b}=\frac{a_{2}}{2\sqrt{\Delta}}+O(\epsilon)\ .\end{split}$ (5)
Here $V_{\rm min}$ is equal to the gain in energy going from the metastable minimum to the origin, while $V_{b}$ is the value of the potential at the barrier.
@@
Another condition to be checked is the tunnelling rate towards the origin for a field configuration starting in the local minimum at high temperature. The action is scale invariant in the limit $T\gg M$, and reads
$\begin{split}\frac{S_{3}(T)}{T}&=4\pi\frac{\tilde{a}_{2}^{3/2}}{\tilde{a}_{3}}B\left(\frac{\tilde{a}_{2}\tilde{a}_{4}}{\tilde{a}_{3}^{2}}\right)\ ,\\ &=4\pi\,\frac{a_{2}^{3/2}}{\Delta}\frac{8\sqrt{2}}{9(\epsilon+2)^{2}(3\epsilon+2)^{3/2}}\,B\left(\frac{2(4+3\epsilon)}{9(2+\epsilon)^{2}}\right)\ ,\end{split}$ (6)
where $\tilde{a}$’s are the coefficients of the potential expanded around the minimum $\phi_{*}(T)$, and they can be expressed in terms of the $a_{i}$’s, with $\tilde{a}_{4}=a_{4}$. Instead, here $B$ is the value of the bounce action for the dimensionless field $\varphi$ in dimensionless three-dimensional euclidean distance $z$,
@@
where $\varphi$ solves the equation of motion with boundary conditions $\varphi^{\prime}(z=0)=0$ with null initial velocity and $\varphi(z\to\infty)=0$, with $B(0)=3.474$ numerically. In our example $\kappa\in[0,2/9)$, and the numerical result is shown in figure 1. We notice two interesting limits of eq. (6). First, when $\epsilon\to 0$ the two minima are degenerate and we are approaching the thin-wall limit, where the action diverges. In particular, we find that $B(2/9-x)\propto x^{-3/2}$ for small $x$. Second, when $\epsilon$ is made arbitrarily large, this corresponds instead to the thick-wall limit, where $B$ approaches a constant. Curiously in this limit the action is suppressed as $\epsilon^{-7/2}$ as $\epsilon$ gets larger and larger. This suppression of the action is due to the lowering of the barrier and its vicinity to the minimum in this limit, as shown in (5). This suppression is contrasted by possible large values of $a_{2}/\sqrt{\Delta}$, which pushes the barrier away from the origin. In both limits, and also away from them, we see that at extremely high-temperature the Euclidean action $S_{3}/T$ is independent from temperature.
@@
Clearly, the thin- and thick-wall cases behave qualitatively different with respect to super-heating. In the thin-wall the action is always quite large, so the tunnelling is rare and the system can really heat up indefinitely in the false vacuum, but the energy available for the phase transition is rather limited, being $\epsilon$-suppressed. In the thick-wall case, the energy can be very large, but the system will like to sample smaller values of the action, leading to faster vacuum decay.
@@ III Superheating from scale invariant sectors
In the previous section we have shown the requirement on the effective potential for $\phi$. Now we identify the microphysics that can lead to such parametric.
@@
where all quartics are positive. We assume that O($N$) symmetry is not spontaneously broken in such a way that when discussing the effective potential for $\phi$, we could always work with $S_{i}=0$. In view of the largeN-N limit that we are going to exploit, we remind that the perturbative regime of our model is attained when the combination
$\bar{\lambda}\equiv\frac{\lambda_{\rm mix}\sqrt{N}\ ,}{16\pi^{2}}$ (9)
is small. Such a condition can be derived by estimating the loop corrections to the $\phi$ quartic coupling, and the same logic can be applied to the quartic coupling for the scalars $S_{i}$, namely $\bar{\lambda}_{S}\equiv\lambda_{S}N/(16\pi^{2})$. See Glioti:2018roy for a very precise derivation and a similar setup.
When the model (8) is put at finite temperature, we can compute the effective potential for $\phi$, and the coefficients $a_{i}$ as they appear in eq. (2). It will be sufficient to work under the assumption that only $S_{i}$ receive thermal corrections at temperature $T$. Therefore we just need to have thermal equilibrium in the system of the $S_{i}$, while we can relax this requirement for $\phi$. This is important in what follows, since $\phi$ can be treated simply as a background field. We notice that our sector can be thought of as a weakly coupled thermal bosonic CFT with a large $N$ number of states, where the coupling to the background field $\phi^{2}$ is the only breaking of scale invariance.
@@
The $J$-function is Boltzmann suppressed at low temperatures, but it is sizable at high temperatures. Let us then focus on the scenario where $T\gg m$, that is let us consider a sector with a temperature always bigger than the mass of the $S_{i}$. In our notation, this corresponds to expand for small $y$. Such an expansion is not analytic at $y=0$ as can be checked by noticing that the coefficient of the term $y^{4}$ is divergent from contributions of small momenta (see for example Laine:2016hma ). The origin of this is at the same time a feature and a worry (see section III.1). Indeed it originates from bosonic zero modes in the three-dimensional theory obtained by compactifying the time direction on a circle of imaginary radius $1/T$. In the three-dimensional theory exists an excitation of mass $m$ (potentially well below the scale $\pi T$) due to periodic boundary conditions on scalar fields, as well as a tower of heavy modes with masses multiples of $2\pi T$.
@@
Let us notice that a contribution linear in $T$ is not present for fermions at finite temperature, since they do not have zero mode excitations in the three-dimensional theory. Both fermions and boson contribute instead in a similar way to $a_{0}$ and $a_{2}$ (and $a_{4}$ with opposite signs), since these terms do not depend much on the contributions from zero modes.
All in all, by expanding at high-temperature the thermal functions for the $S_{i}$, we get the following contributions to $\phi$ potential of eq. (2)
@@
Let us comment on the size of these coefficients at largeN-N. Clearly, the thermal pressure of the particles is enhanced by $N$. The other coefficients, instead, show a different behaviour with $N$ when expressed in terms of $\bar{\lambda}$, which we hold fixed and perturbative. In particular, we have that $a_{2}\propto\bar{\lambda}\sqrt{N}$ and $a_{3}\propto 1/\sqrt{N}$. This shows that we can have parametrically large values of $a_{2}$ while still being perturbative. Also, since given the scaling of $a_{3}$ we cannot really take the limit $N\to\infty$, as this imposes an upper bound.
In the model under consideration, the conditions in Eqs. (3) and (4) are independent of $N$, yet together they imply the constraint $2\sqrt{2\lambda_{0}}/3<16\pi\bar{\lambda}<\sqrt{\lambda_{0}}$. At the same time, requiring the barrier lie sufficiently far from the origin—so that re-summation can be avoided (cf. the next subsection)—yields $16\pi^{2}\bar{\lambda}\lesssim\sqrt{N}/9$. Furthermore, since the bounce action scales as $S_{3}/T\propto N^{1/4}$, there is an upper bound on $N$ to ensure the tunnelling condition is satisfied. This, as exemplified by figure 2, leads to $\bar{\lambda}\approx 0.015$ and $N\approx 250$.
@@
This discussion has a very strong analogy with the one of Ref.  Glioti:2018roy , where the authors have identified a mechanism to break electroweak symmetry at arbitrarily high temperatures. In analogy, here we still want a symmetry-breaking minimum at high temperatures, but differently from that work, we are after a metastable minimum leading to a first order phase transition. This leads us to work in a limit where resummation is not important as not to suppress the contribution to the barrier as we discuss next.

### v2

@@ I Introduction
The temperature of the universe has been decreasing ever since the end of inflation, when the Standard Model (SM) was reheated at its highest temperatures. Throughout the standard evolution of our Universe, the temperature has been decreasing relentlessly, possibly crossing the critical values, $T_{c}$, of some phase transitions. If these phase transitions were first order, there would be a potentially detectable signal in the form of a stochastic gravitational waves (GWs) background with a characteristic peak frequency, see Caprini and others (2016); Hindmarsh et al. (2014, 2015, 2017); Cutting et al. (2018, 2020); Gould et al. (2021); Weir (2018); Hindmarsh et al. (2021); Caprini and others (2019); Badger and others (2023); Caldwell and others (2022). Unfortunately, in the SM, neither the electroweak symmetry breaking nor QCD confinement are first order phase transitions. Nonetheless, these are targets of current interferometers Abbott and others (2019) and PTAs Agazie and others (2023); Antoniadis and others (2023) as well as many future observatories such as LISA Amaro-Seoane and others (2017) and ET Sathyaprakash and others (2012). This is due to the fact that it is conceivable that new physics can, for example, affect the shape of the Higgs potential, making the electroweak symmetry breaking a first order phase transition Grojean et al. (2005); Delaunay et al. (2008); Ellis et al. (2018, 2019); Bruggisser et al. (2018); Espinosa and Quiros (2007); Beniwal et al. (2017); Barger et al. (2008); Espinosa et al. (2012); Kozaczuk et al. (2020); Kurup and Perelstein (2017); Azatov et al. (2022) (see Athron et al. (2024) for a review). More generally, other phase transitions can be present associated with the breaking of some symmetries, each with clear experimental targets at GW interferometers.
@@
As emphasised in Buen-Abad et al. (2023), however, studying epochs where the temperature increases - while the Universe expands little - can reveal interesting new effects. Indeed, if the temperature grows above $T_{c}$ and then back below it, it might well be the case that a phase transition happens twice.
In this work, we explore this idea further, trying to understand what are the particle physics models that can show a phase transition when their sectors are heating up, possibly thanks to an external reservoir. In most cases we will focus on the reheating phase of the Universe, but our results will apply to more generic scenarios (such as a freeze-in type of generation). Also, we want to identify which are those models that can be heated up maximally, in the search for large signals in GWs.
We identify scenarios of super-heating in sectors with a large number of light bosons in thermal equilibrium, as they have the properties to generate a persistent meta-stable minimum where these bosons are actually massive. The order parameter is a generic background field $\phi$, which, rather interestingly instead, does not need to be thermal. In these sectors, there is no a priori upper limit on the maximal reheating temperature, but only a practical limitation induced by the time scale for nucleation tunnelling. [^1] The term “tunnelling” here is used to denote the finite-temperature decay of a metastable vacuum described by an $O(3)$-symmetric Euclidean bounce with action $S_{3}/T$, following the terminology originally introduced in Linde (1983). The first order transition then proceeds with bubble nucleation towards the new minimum where the states are massless. The interplay between temperature and the mass difference results in transitions of inverse type Barni et al. (2024, 2025), indeed, the plasma particles gain momentum in the true vacuum where they are relativistic Buen-Abad et al. (2023). Other setups of phase transitions while heating can be found in Caprini and No (2012); Casalderrey-Solana et al. (2022); Kolesova and Laine (2024); Dent et al. (2024); Ai et al. (2024).
We assess the importance of two aspects underlying super-heating and heating transitions in general. First, the sector has to have a very largeN-N number of bosons in order to populate parameter space where the effective potential for $\phi$ is under perturbative control, and the calculation of the nucleation rate is reliable. Second, the sector has to be heated-up slowly, so as to ensure that the zero temperature minimum, where the bosons are massive, becomes a metastable minimum also at high temperature. We find that efficient thermal contact with the external reservoir at temperatures below the critical value during the heating phase is essential. In this regime, the system can undergo a pair of phase transitions—an inverse transition upon heating followed by a direct transition upon cooling—and we quantify the resulting characteristic double-peaked gravitational-wave signal.
@@ II Super-heating: generalities
In this section, we sketch the basic idea behind a superheating phase transition. The aim is to highlight the needed ingredients to have a potential that has a metastable minimum that persists up to extremely high temperatures. We are mostly interested in situations where the metastability of this configuration leads to a first order phase transition via nucleation of bubbles, as this has associated production of gravitational waves from the early universe, but we will comment on more generic endpoints.
Let us consider the example of a sector with a scalar quantum field, $\phi$, at finite temperature, $T$. As it will be evident, the sector needs additional interacting light fields, but for the moment, we do not discuss the microphysics of this setup. The typical mass scale of $\phi$ is $M$, such that at zero temperature the field has potential
@@
For $M^{2}>0$ the field breaks spontaneously a $Z_{2}$ symmetry, since $\phi=0$ is a maximum, in a minimum $\langle\phi\rangle=f$. [^2] Notice also that we are just interested in classifying the properties of super-heating, the fact of using a real scalar is just for simplicity. Therefore, we are not going to discuss the domain wall generation when $Z_{2}$ is spontaneously broken. Eq. (1) is just prototypical of a sector with spontaneous symmetry breaking.
@@
The different normalizations in Eqs. (1) and (2), reflect distinct conventions for the tree-level and finite-temperature potentials, and are related by simple redefinitions of the coefficients. If the temperature is the largest scale, and we can neglect $M$, this is also all the potential for $\phi$, and we will discuss later in section IV the impact of finite $M$. In conventional field theories, we have $a_{1}=0$, $a_{3}\leq 0$ and $a_{4}>0$, while the sign of $a_{2}$ depends on the theory Weinberg (1974), but it is almost always positive. Also, usually, the field is subject to the constraint $\phi\geq 0$.  [^3] Restricting the analysis to $\phi\geq 0$ is a matter of parametrization and does not entail any loss of generality for a quartic potential bounded from below. In fig. 1 we show the effective potential at different temperatures. At low temperature the minimum away from the origin is the global one, while at higher temperatures the configuration at the origin becomes energetically favoured. The superheating regime persists as long as the potential barrier separating the two minima is present at high temperatures.
Figure 1: Finite–temperature effective potential $V(\phi,T)$, normalised to $|V(\phi_{\ast},T_{0})|$, shown as a function of $\phi/\phi_{\ast}(T_{0})$ at the characteristic temperatures $T_{0}$ (blue), $T_{c}$ (black dashed), and $T_{\rm SP}$ (red) (see sec. IV). The plot illustrates the presence of a metastable minimum away from the origin, its degeneracy with the symmetric phase at $T_{c}$, and its disappearance at the spinodal temperature $T_{\rm SP}$. The orange and green curves correspond to intermediate nucleation temperatures, $T_{n}^{\leftarrow}$ and $T_{n}^{\rightarrow}$ respectively (see sec. IV.5), while arrows indicate the inverse (orange) and direct (green) transition paths. Dotted light red curves denote the potential for higher temperatures in the case when superheating does not last forever.
@@
where the second ensures that the new minimum - which we call $\phi_{*}(T)$ from now on - is in the region $\phi\geq 0$.
Notice that crucially, in this regime of ultra-high temperature, all these conditions are temperature independent. This is due to the fact that eq. (2) can be rewritten in terms of a dimensionless combination $\varphi=\phi/T$, and the potential will just show an overall $T^{4}$ factor.
@@
Let us emphasise that if both (3) and (4) are satisfied, then the potential (2) has a metastable minimum away from the origin for any value of the temperature, including extremely high values. We are therefore interested in a slightly different phase than the one mainly discussed in Buen-Abad et al. (2023).
If interested in a local minimum $\phi_{*}(T)$ at high temperature, it is convenient to consider models where $a_{2}$ is very large, such as to push the barrier, located at $\phi_{b}(T)$, and the local minimum away from the origin, to improve calculability. Notice that the barrier and the local minimum are always separated by $\phi_{*}(T)-\phi_{b}(T)=(\sqrt{\Delta}/a_{4})\,T$, located at an average distance from the origin $\approx-a_{3}/(2a_{4})T$ for small $\Delta$. Taking $\Delta$ as an input, we can express $a_{4}=(a_{3}^{2}-\Delta)/4a_{2}$, which leads to $a_{3}<-3\sqrt{\Delta}$ as an implicit new form for eq. (4). For this, it is convenient to define $a_{3}=-3(1+\epsilon)\sqrt{\Delta}$, with $\epsilon>0$. Therefore we have
$\begin{split}V_{\rm min}&=T^{4}\frac{a_{2}^{3}\epsilon}{\Delta(2+3\epsilon)^{3}}\ ,\\ \frac{V_{b}}{V_{\rm min}}&=\frac{1}{4\epsilon}+O(\epsilon)\ ,\quad\varphi_{b}=\frac{a_{2}}{2\sqrt{\Delta}}+O(\epsilon)\ ,\end{split}$ (5)
where we recall that $\varphi_{b}=\phi_{b}/T$ and $V_{b}=V(\phi_{b})$. Here $V_{\rm min}$ is equal to the gain in energy going from the metastable minimum to the origin, while $V_{b}$ is the value of the potential at the barrier.
@@
Another condition to be checked is the nucleation rate towards the origin for a field configuration starting in the local minimum at high temperature. The action is scale invariant in the limit $T\gg M$, and reads
$\begin{split}\frac{S_{3}(T)}{T}&=4\pi\frac{\tilde{a}_{2}^{3/2}}{\tilde{a}_{3}}B\left(\frac{\tilde{a}_{2}\tilde{a}_{4}}{\tilde{a}_{3}^{2}}\right)\ ,\\ &=4\pi\,\frac{a_{2}^{3/2}}{\Delta}\frac{8\sqrt{2}}{9(\epsilon+2)^{2}(3\epsilon+2)^{3/2}}\,B\left(\frac{2(4+3\epsilon)}{9(2+\epsilon)^{2}}\right)\ .\end{split}$ (6)
This expression follows from the standard result for the three–dimensional Euclidean action of a quartic potential, $S_{3}/T=4\pi\,(a_{2}^{3/2}/a_{3})\,B(a_{2}a_{4}/a_{3}^{2}),$ which is usually derived when studying nucleation from the origin toward a non-zero vacuum expectation value, see Sarid (1998). In the present case, however, we are interested in the inverse process, namely the nucleation from the metastable minimum at $\phi=\phi_{*}(T)$ back to the origin. This is conveniently described by expanding the finite-temperature effective potential around the local minimum $\phi_{*}(T)$ and identifying the coefficients of the resulting quartic expansion. The quantities $\tilde{a}_{i}$ denote these expansion coefficients, which can be expressed in terms of the original parameters $a_{i}$ of the potential, with $\tilde{a}_{4}=a_{4}$. Instead, here $B$ is the value of the bounce action for the dimensionless field $\varphi$ in dimensionless three-dimensional Euclidean distance $z$,
@@
where $\varphi$ solves the equation of motion with boundary conditions $\varphi^{\prime}(z=0)=0$ with null initial velocity and $\varphi(z\to\infty)=0$, with $B(0)=3.474$ numerically. In our example $\kappa\in[0,2/9)$, and the numerical result is shown in figure 2. We notice two interesting limits of eq. (6). First, when $\epsilon\to 0$ the two minima are degenerate and we are approaching the thin-wall limit, where the action diverges. In particular, we find that $B(2/9-x)\propto x^{-3/2}$ for small $x$. Second, when $\epsilon$ is made arbitrarily large, this corresponds instead to the thick-wall limit, where $B$ approaches a constant. Curiously in this limit the action is suppressed as $\epsilon^{-7/2}$ as $\epsilon$ gets larger and larger. This suppression of the action is due to the lowering of the barrier and its vicinity to the minimum in this limit, as shown in (5). This suppression is contrasted by possible large values of $a_{2}/\sqrt{\Delta}$, which push the barrier away from the origin. In both limits, and also away from them, we see that at extremely high-temperature the Euclidean action $S_{3}/T$ is independent of temperature.
@@
Clearly, the thin- and thick-wall cases behave qualitatively differently with respect to super-heating (see fig. 1), which illustrates how these two regimes correspond to markedly distinct shapes of the finite-temperature potential: in the thin-wall limit the two minima are nearly degenerate and separated by a high and narrow barrier, while in the thick-wall regime the barrier is lower and broader, bringing the metastable minimum closer to instability. In the thin-wall the action is always quite large, so the nucleation is rare and the system can really heat up indefinitely in the false vacuum, but the energy available for the phase transition is rather limited, being $\epsilon$-suppressed. In the thick-wall case, the energy can be very large, but the system will like to sample smaller values of the action, leading to faster vacuum decay.
@@ III Superheating from scale invariant sectors
In the previous section we have shown the requirement on the effective potential for $\phi$. Now we identify the microphysics that can lead to this parametric behaviour.
@@
where all quartics are positive. We assume that O($N$) symmetry is not spontaneously broken in such a way that when discussing the effective potential for $\phi$, we can always work with $S_{i}=0$. In view of the largeN-N limit that we are going to exploit, we remind that the perturbative regime of our model is attained when the combination
$\bar{\lambda}\equiv\frac{\lambda_{\rm mix}\sqrt{N}}{16\pi^{2}}\ ,$ (9)
is small. Such a condition can be derived by estimating the loop corrections to the $\phi$ quartic coupling, and the same logic can be applied to the quartic coupling for the scalars $S_{i}$, namely $\bar{\lambda}_{S}\equiv\lambda_{S}N/(16\pi^{2})$. See Glioti et al. (2019) for a very precise derivation and a similar setup.
When the model (8) is placed at finite temperature, we can compute the effective potential for $\phi$, and the coefficients $a_{i}$ as they appear in eq. (2). It will be sufficient to work under the assumption that only $S_{i}$ receive thermal corrections at temperature $T$. Therefore we just need to have thermal equilibrium in the system of the $S_{i}$, while we can relax this requirement for $\phi$. This is important in what follows, since $\phi$ can be treated simply as a background field. We notice that our sector can be thought of as a weakly coupled thermal bosonic CFT with a large $N$ number of states, where the coupling to the background field $\phi^{2}$ is the only breaking of scale invariance.
@@
The $J$-function is Boltzmann suppressed at low temperatures, but it is sizable at high temperatures. Let us then focus on the scenario where $T\gg m$, that is let us consider a sector with a temperature always bigger than the mass of the $S_{i}$. In our notation, this corresponds to expand for small $y$. Such an expansion is not analytic at $y=0$ as can be checked by noticing that the coefficient of the term $y^{4}$ is divergent from contributions of small momenta (see for example Laine and Vuorinen (2016)). The origin of this is at the same time a feature and a worry (see section III.1). Indeed it originates from bosonic zero modes in the three-dimensional theory obtained by compactifying the time direction on a circle of imaginary radius $1/T$. In the three-dimensional theory exists an excitation of mass $m$ (potentially well below the scale $\pi T$) due to periodic boundary conditions on scalar fields, as well as a tower of heavy modes with masses multiples of $2\pi T$.
@@
Let us notice that a contribution linear in $T$ is not present for fermions at finite temperature, since they do not have zero mode excitations in the three-dimensional theory. Both fermions and bosons contribute instead in a similar way to $a_{0}$ and $a_{2}$ (and $a_{4}$ with opposite signs), since these terms do not depend much on the contributions from zero modes.
All in all, by expanding at high-temperature the thermal functions for the $S_{i}$, we get the following contributions to the $\phi$ potential of eq. (2)
@@
Let us comment on the size of these coefficients at largeN-N. Clearly, the thermal pressure of the particles is enhanced by $N$. By contrast, the other coefficients show a different behaviour with $N$ when expressed in terms of $\bar{\lambda}$, which we hold fixed and perturbative. In particular, we have that $a_{2}\propto\bar{\lambda}\sqrt{N}$ and $a_{3}\propto 1/\sqrt{N}$. This shows that we can have parametrically large values of $a_{2}$ while still being perturbative. Also, since given the scaling of $a_{3}$ we cannot really take the limit $N\to\infty$, as this imposes an upper bound.
In the model under consideration, the conditions in Eqs. (3) and (4) are independent of $N$, yet together they imply the constraint $2\sqrt{2\lambda_{0}}/3<16\pi\bar{\lambda}<\sqrt{\lambda_{0}}$. At the same time, requiring the barrier lie sufficiently far from the origin—so that re-summation can be avoided (cf. the next subsection)—yields $16\pi^{2}\bar{\lambda}\lesssim\sqrt{N}/9$. Furthermore, since the bounce action scales as $S_{3}/T\propto N^{1/4}$, there is an upper bound on $N$ to ensure the nucleation condition is satisfied. This, as exemplified by figure 3, leads to $\bar{\lambda}\approx 0.015$ and $N\approx 250$.
@@
This discussion has a very strong analogy with the one of Ref.  Glioti et al. (2019), where the authors have identified a mechanism to break electroweak symmetry at arbitrarily high temperatures. In analogy, here we still want a symmetry-breaking minimum at high temperatures, but differently from that work, we are after a metastable minimum leading to a first order phase transition. This leads us to work in a limit where resummation is not important as not to suppress the contribution to the barrier as we discuss next.